OpenStudy (anonymous):

If n and p are integers greater than 1 5n is the square of a number 75np is the cube of a number. The smallest value for n + p is? A. 14 B. 18 C. 20 D. 30 E. 50

OpenStudy (anonymous):

$$n, p > 1 \\ 5n = x^2\\ 75np = x^3$$ Solve for n in terms of x: $$n = \frac {x^2} 5$$ $$75(\frac{x^2} 5)p = x^3$$ Solve for p in terms of x: $$15x^2p = x^3 \\ p = 15x$$ $$\frac{x^2} 5 + 15x = m$$ We wish to minimize \(m\).

OpenStudy (anonymous):

C. 20

OpenStudy (anonymous):

@bloopman your method is invalid

OpenStudy (anonymous):

@TanteAnna, no, it isn't. I was about to use calculus to solve it, but because you were so kind to give him the answer I decided not to. Perhaps you should read the ToS before you give out another answer like that.

OpenStudy (anonymous):

Ahaha! When you think that 5 and 75 are the same number your method is certainly invalid! Quit calling me out when you know you are wrong! How about you go read the ToS! @bloopman

OpenStudy (anonymous):

No need to be aggressive. Please, do tell me where I stated 5 = 75.

OpenStudy (anonymous):

5n=x275np=x3

OpenStudy (anonymous):

That's where.

OpenStudy (anonymous):

How is that saying 5 = 75?

OpenStudy (anonymous):

Because you say x is the number in both equations but in the first equation we want 5 and in the second equation we want 75 so you are implying that 5 = x = 75 therefore 5 = 75. don't be a noob next time thanks

OpenStudy (anonymous):

...

OpenStudy (anonymous):

Lol'd. 10/10 if troll 7/10 if copypasta

OpenStudy (anonymous):

You are the only troll here